# Spatial-Aware Decision-Making with Ring Attractors in Reinforcement Learning Systems

### v2

@@ 5 Conclusion
This paper presents a novel approach to RL, integrating ring attractors into action selection. Our work demonstrates that these neuroscience-inspired ring attractors significantly enhance learning capabilities for value-based RL agents, leading to more stable and efficient action selection, particularly in spatially structured tasks.
@@
The integration of ring attractors as a DL module proves to be particularly effective, allowing for end-to-end training and easy incorporation into existing RL architectures. Our experiments demonstrate significant improvements in action selection and learning speed, achieving state-of-the-art results on the Atari 100k benchmark. Notably, we observe an average 53% performance increase across all tested games compared to prior baselines. The most relevant improvements were observed in games with strong spatial components, such as Asterix (110 % improvement) and Boxing (105% improvement). Additionally, we observed improvements in other environments tested outside the Atari benchmark, further supporting the effectiveness of our approach across various RL tasks and agents.
6 Future research
@@
Future research should investigate scalability in high-dimensional action spaces and explore its efficacy in domains where spatial relationships are less straightforward. We believe that the success of this approach opens up several future research paths. The current work can be extended to multi-agent scenarios and policy-based RL agents. In the field of uncertainty-aware decision making, leveraging the spatial structure provided by attractor networks presents a promising avenue to map uncertainty explicitly to the action space. Deploying the techniques presented here into specific domains could yield performance boosts, especially in safe RL, leveraging their stability properties to enforce constraints and ensure predictable behavior. This approach not only improves performance, but also offers potential insight into spatial encoding of actions and decision-making processes, bridging the gap between neuroscience-inspired models and practical RL agents.
@@ A.1 Background:Attractor Networks Theoretical Foundations
Ring attractor networks are a type of biological neural structure that has been proposed to underlie the representation of various cognitive functions, including spatial navigation, working memory, and decision-making (Kim et al., 2017).
@@
Biological intuition. In the early 1990s, the research carried out by  Zhang (1996) proposed that ring neural structures could underlie the representation of heading direction in rodents. Zhang (1996) argued that the neural activity in an attractor network might encode the direction of the animal’s head, with the network transitioning from one attractor state to another state as the animal turns.
@@
Empirical evidence. There is growing evidence from neuroscience supporting the role of ring attractors in neural processing. For example, electrophysiological recordings from head direction cells (HDCs) of rodents have revealed a circular organisation of these neurons, with neighbouring HDCs encoding slightly different heading directions (Taube, 1995). Furthermore, studies have shown that HDC activity can be influenced by sensory inputs, such as visual signals and vestibular signals, and that these inputs can cause the network to update its representation of heading direction (Taube, 2007).  Xiong & Sridhar (2024) showed model internal noise biases toward accuracy over speed, while environmental uncertainty exhibits a U-shaped effect where moderate uncertainty favors speed and extreme uncertainty favors accuracy.  Wilson (2023) demonstrated that biological navigation networks rely on attractor dynamics while maintaining adaptability through continuous synaptic plasticity and sensory feedback.
@@
Sensor fusion in ring attractors. Ring attractor networks provide a theoretical foundation for understanding cognitive functions such as spatial navigation, working memory, and decision-making. In the context of action selection in RL, sensor fusion plays a pivotal role in augmenting the information-processing capabilities of these networks. By combining data from various sensory modalities, ring attractors create a more nuanced and robust representation of the environment, essential for adaptive behaviors (ME, 2008). Research has elucidated the relationship between ring attractors and sensory inputs, with the circular organisation of HDCs in rodents complemented by the convergence of visual and vestibular inputs, highlighting the integrative nature of sensory information within the ring attractor framework (Zugaro et al., 2001).
@@
Modulation by sensory inputs. Beyond the spatial domain, sensory input dynamically influences the activity of ring attractor networks. Studies have shown that visual cues and vestibular signals not only update the representation of heading direction but also contribute to the stability of attractor states, allowing robust spatial memory and navigation (Goodridge et al., 1998).
@@
Sensor fusion for action selection. The concept of sensor fusion within the context of ring attractors extends beyond traditional sensory modalities, encompassing diverse sources, such as proprioceptive and contextual cues (McNaughton et al., 1996). Building on the foundation of ring attractor networks discussed earlier, the integration of sensor fusion in the context of action selection involves fusing the action values associated with each potential action within the ring attractor framework. In particular, sensory information, previously shown to modulate the activity of ring attractor networks, extends its influence to the representation of action values. The inclusion of sensory information reflects a higher cognitive process, where the adaptable nature of ring attractor networks plays a central role in orchestrating optimal decision making and action selection in complex environments.
A.2 Validating Ring Attractor Contributions Through Ablation Studies
A.2.1 Exogenous Ring Attractor Model Ablation Study
@@
Figure 4: Ablation study comparing BDQN variants in OpenAI Gym Super Mario Bros (Kauten, 2018). The plot shows cummulative reward over 1 million frames for three models: Standard BDQN (Azizzadenesheli et al., 2018) ; BDQNRA-UA with RA and Uncertainty Awareness (UA) implementing both the ring attractor behavior policy from Section 3.1.2 and the uncertainty quantification model from 3.1.3; and BDQNRA-RM, applying the same concepts from BDQNRA-UA, but randomly distributing the action space across the ring in each experiment. Displaying mean episodic returns over 10 averaged seeds.
A.2.2 Deep Learning Ring Attractor Model Ablation Study
@@
Figure 5: Ablation study results comparing the performance of the full RNN-based ring attractor model against a version with the circular weight distribution removed. The graph illustrates a significant performance drop for the Ms Pacman and Chopper Command environments in the Atari 100K benchmark (Bellemare et al., 2012). This emphasises the role of the circular topology in encoding spatial information and enhancing learning. Displaying mean episodic returns over 10 averaged seeds.
@@
A.2.3 Deep Learning Ring Attractor Model Evolution
In this appendix section, we analyse model dynamics with both forward pass ($V(s)$) and hidden-to-hidden ($U(v)$) weights made trainable, rather than the standard approach of fixed forward pass connections, as presented in Section 3.2. As shown in Fig. 6, the forward pass connections preserve the ring structure over training time, with strong distance-dependent decay patterns maintained throughout the learning process. This may indicate that the network naturally favors maintaining spatial topology for transmitting sensory information on a per-frame basis.
@@ A.2.3 Deep Learning Ring Attractor Model Evolution
The hidden-to-hidden connections, depicted in Fig. 7, demonstrate markedly different behavior. These connections evolve beyond their initial ring structure, developing specialized patterns that enable the encoding of environment-specific relationships between neurons in the hidden space. This flexibility in hidden layer connectivity supports the learning of complex action relationships while building upon the structured spatial representation from the forward pass.
These findings validate our standard implementation approach described in Section 3.2, where forward pass connections are fixed and only hidden-to-hidden weights remain trainable. The natural preservation of ring structure in trainable forward weights may suggest this topology is inherently beneficial for processing spatial information, while adaptable hidden weights enable the task-specific learning demonstrated in our experimental results, Section 4.2.

### v3

@@ Conclusion.
This paper presents a novel approach to RL, integrating ring attractors into action selection. Our work demonstrates that these neuroscience-inspired ring attractors significantly enhance learning capabilities for RL agents, leading to more stable and efficient action selection, particularly in spatially structured tasks.
@@
The integration of ring attractors as a DL module proves to be particularly effective, allowing for end-to-end training and easy incorporation into existing RL architectures. Our experiments demonstrate significant improvements in action selection and learning speed, achieving state-of-the-art results on the Atari 100K benchmark. Notably, we observe an average 53% performance increase across all tested games compared to prior baselines. The most relevant improvements were observed in games with strong spatial components, such as Asterix (110 % improvement) and Boxing (105% improvement). Additionally, we observed improvements in other environments tested outside the Atari benchmark, further supporting the effectiveness of our approach across RL tasks and agents.
Future work.
Future research should investigate scalability in high-dimensional action spaces and explore its efficacy in domains where spatial relationships are less straightforward. We believe that the success of this approach opens up several future research paths. This work can be extended to multi-agent scenarios and continuous control tasks as these have been the current limits of this piece of research. The ring attractor approach may exhibit reduced applicability in discrete action spaces with cardinality $|\mathcal{A}|<3$ where the circular topology overhead exceeds potential spatial encoding benefits, and in action spaces where the action correlation matrix approaches the identity matrix, indicating absence of spatial dependencies or sequential relationships between discrete actions. Additionally, the extension to 3D spatial navigation tasks presents an organic expansion of this work, these scenarios will naturally introduce realistic sensor noise into the system.
@@
In the field of uncertainty-aware decision making, leveraging the spatial structure provided by attractor networks presents a promising avenue to map uncertainty explicitly to the action space. Deploying the techniques presented here into specific domains could yield performance boosts, especially in safe RL, leveraging their stability properties to enforce constraints and ensure predictable behavior. This approach not only improves performance but also offers potential insight into spatial encoding of actions and decision-making processes, bridging the gap between neuroscience-inspired models and practical RL agents.
Acknowledgments and Disclosure of Funding
This work was conducted as part of a self-funded PhD programme at the University of Manchester. Computational resources were provided by the Computational Shared Facility 3 (CSF3) at the University of Manchester csf3 . We gratefully acknowledge the support of the Research IT team for maintaining and providing access to the GPU Cluster.
Funding: This research received no external funding.
Competing Interests: No competing interests are declared.
@@ A.1 Background: Attractor Networks Theoretical Foundations
Ring attractor networks are a type of biological neural structure that has been proposed to underlie the representation of various cognitive functions, including spatial navigation, working memory, and decision-making [18].
@@
Biological intuition. In the early 1990s, the research carried out by [50] proposed that ring neural structures could underlie the representation of heading direction in rodents. [50] argued that the neural activity in an attractor network might encode the direction of the animal’s head, with the network transitioning from one attractor state to another state as the animal turns.
@@
Empirical evidence. There is growing evidence from neuroscience supporting the role of ring attractors in neural processing. For example, electrophysiological recordings from head direction cells (HDCs) of rodents have revealed a circular organization of these neurons, with neighbouring HDCs encoding slightly different heading directions [38]. Furthermore, studies have shown that HDC activity can be influenced by sensory inputs, such as visual signals and vestibular signals, and that these inputs can cause the network to update its representation of heading direction [37][47] showed model internal noise biases toward accuracy over speed, while environmental uncertainty exhibits a U-shaped effect, where moderate uncertainty favors speed and extreme uncertainty favors accuracy. [46] demonstrated that biological navigation networks rely on attractor dynamics while maintaining adaptability through continuous synaptic plasticity and sensory feedback.
@@
Sensor fusion in ring attractors. Ring attractor networks provide a theoretical foundation for understanding cognitive functions such as spatial navigation, working memory, and decision-making. In the context of action selection in RL, sensor fusion plays a pivotal role in augmenting the information-processing capabilities of these networks. By combining data from various sensory modalities, ring attractors create a more nuanced and robust representation of the environment, essential for adaptive behaviors [24]. Research has elucidated the relationship between ring attractors and sensory inputs, with the circular organisation of HDCs in rodents complemented by the convergence of visual and vestibular inputs, highlighting the integrative nature of sensory information within the ring attractor framework [51].
@@
Modulation by sensory inputs. Beyond the spatial domain, sensory input dynamically influences the activity of ring attractor networks. Studies have shown that visual cues and vestibular signals not only update the representation of heading direction but also contribute to the stability of attractor states, allowing robust spatial memory and navigation [12].
@@
Sensor fusion for action selection. The concept of sensor fusion within the context of ring attractors extends beyond traditional sensory modalities, encompassing diverse sources, such as proprioceptive and contextual cues [23]. Building on the foundation of ring attractor networks discussed earlier, the integration of sensor fusion in the context of action selection involves fusing the action values associated with each potential action within the ring attractor framework. In particular, sensory information, previously shown to modulate the activity of ring attractor networks, extends its influence to the representation of action values. The inclusion of sensory information reflects a higher cognitive process, where the adaptable nature of ring attractor networks supports optimal decision-making and action selection in complex environments.
A.2 Function Approximation for Q-Values
A.2.1 Motivation and Background
In RL, an agent optimizes behaviour by interacting with an environment modeled as a Markov Decision Process (MDP), $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is the state transition probability function, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function, and $\gamma\in[0,1)$ denotes the discount factor. The agent seeks to learn a policy $\pi(a\mid s)$, which defines the probability of selecting action $a$ in state $s$, that maximizes the expected discounted return [34]:
$J(\pi)=\mathbb{E}_{\tau}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right],$ (13)
where $J(\pi)$ is the objective function representing the expected total return (performance measure) of policy $\pi$, $\tau=(s_{0},a_{0},r_{0},s_{1},a_{1},r_{1},\ldots)$ represents a trajectory of states, actions, and rewards, $r_{t}$ is the reward received at time step $t$, $\gamma$ is the discount factor that balances immediate versus future rewards, and the expectation is taken over trajectories sampled by following policy $\pi$.
Action-value functions are foundational to various value-based RL algorithms. The action-value function is defined as:
$Q^{\pi}(s,a)=\mathbb{E}_{\tau}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\;\middle|\;s_{0}=s,\,a_{0}=a,\,\pi\right],$ (14)
which quantifies the expected cumulative discounted reward when executing action $a$ in state $s$ and subsequently following policy $\pi$ for all future time steps.
This function satisfies a recursive relationship known as the Bellman equation [6]:
$Q^{\pi}(s,a)=\mathbb{E}_{s^{\prime},r\sim P,R}\left[r+\gamma\sum_{a^{\prime}}\pi(a^{\prime}\mid s^{\prime})Q^{\pi}(s^{\prime},a^{\prime})\right],$ (15)
where $\mathbb{E}_{s^{\prime},r\sim P,R}$ denotes expectation with respect to the next state $s^{\prime}$ and reward $r$ following the environment dynamics defined by $P$ and $R$, and $r+\gamma\sum_{a^{\prime}}\pi(a^{\prime}\mid s^{\prime})Q^{\pi}(s^{\prime},a^{\prime})$ is the expected immediate reward plus the discounted value of future actions. This forms the mathematical basis for value iteration, policy iteration, Q-learning, and other dynamic programming and temporal-difference methods used to compute or approximate $Q^{\pi}$.
In environments with large or continuous state-action spaces, storing tabular representations of $Q(s,a)$ becomes infeasible due to memory constraints and the inability to visit all state-action pairs sufficiently often. Even in moderately sized discrete domains, function approximation is often employed to leverage generalization across similar states or actions, thereby speeding up learning and reducing data requirements. By using function approximators, RL algorithms can handle high-dimensional inputs and share learned structure across different parts of the state-action space.
A.2.2 General Formulation
Q-functions are approximated using parameterized functions $Q(s,a;\theta)$, where $\theta\in\mathbb{R}^{d}$ are learnable weights:
$Q(s,a;\theta)\approx Q^{\pi}(s,a),$ (16)
where the target $Q^{\pi}$ could be the Q-function of the current policy (policy evaluation) or the optimal Q-function $Q^{*}$ (control). The goal is to learn $\theta$ such that $Q(s,a;\theta)$ closely estimates the target Q-values across the relevant state-action space.
Parameters $\theta$ are typically updated through stochastic gradient descent to minimize a loss function. For instance, using the mean squared Bellman error:
$\mathcal{L}(\theta)=\mathbb{E}_{s,a,r,s^{\prime}}\left[\bigl(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime};\theta^{\prime})-Q(s,a;\theta)\bigr)^{2}\right],$ (17)
where $\theta^{\prime}$ represents parameters of a target network that is periodically updated to match $\theta$ (in deep Q-learning algorithms such as DQN [26], or $\theta^{\prime}=\theta$ in simpler settings like traditional Q-learning). The term $r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime};\theta^{\prime})$ is often called the “target” value, while $Q(s,a;\theta)$ is the “prediction”. The parameters are updated according to:
$\theta_{t+1}=\theta_{t}-\alpha\nabla_{\theta}\mathcal{L}(\theta_{t}),$ (18)
where $\alpha$ is the learning rate, and $\nabla_{\theta}\mathcal{L}(\theta_{t})$ is the gradient of the loss function with respect to $\theta_{t}$. In practice, these updates are performed using stochastic gradient descent or variants on mini-batches of experience to improve computational efficiency and stability.
A.2.3 Linear Function Approximation
A straightforward approach is linear function approximation, where state-action pairs are mapped into a feature space via a feature function $\phi:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}^{d}$, and Q-values are computed as:
$Q(s,a;\theta)=\theta^{\top}\phi(s,a).$ (19)
This formulation is computationally efficient and, under specific conditions, enjoys theoretical convergence guarantees in on-policy settings [35]. However, linear approximation’s representational capacity is fundamentally constrained by the chosen feature representation $\phi(s,a)$. For discrete action spaces, an alternative formulation represents a separate linear function for each action:
$Q(s,a;\theta)=\theta_{a}^{\top}\phi(s),$ (20)
where $\phi(s)\in\mathbb{R}^{d}$ is a state feature vector and $\theta_{a}\in\mathbb{R}^{d}$ is a parameter vector specific to action $a$.
The transition from linear to neural approximation typically retains this foundational structure while enhancing the feature extraction process. Instead of relying on handcrafted features, neural networks learn feature representations automatically through their hidden layers.
A.2.4 Neural Function Approximators
To capture richer, nonlinear dependencies, neural networks can be used to parameterize Q-functions. In the field of DRL, for any state-action pair $(s,a)$, the Q-value $Q(s,a)$ can be expressed as a function of the input state through a function approximation algorithm $\Phi_{\theta}(s)$. This notation evolves from the linear case but generalizes to allow for arbitrary nonlinear transformations of the input. Formally:
$Q(s,a)=\Phi_{\theta}(s)=\theta^{\top}x(s),$ (21)
where $x(s)$ corresponds to the (learned) feature representation of the state $s$ (e.g., the outputs of the penultimate layer), and $\theta$ represents the weights of the final layer.
With this design, neural networks learn both the feature extraction ($x(s)$) and the final mapping ($\theta$) end-to-end through backpropagation, enabling highly expressive function approximators that can generalize across large or complex state-action domains.
A.3 Policy Gradient Methods with Ring Attractors
While previous sections focused on value-based methods, ring attractors can also enhance policy gradient methods, such as Proximal Policy Optimization (PPO) [29]. The integration follows similar principles, but adapts to the unique characteristics of policy-based approaches [36].
A.3.1 Proximal Policy Optimization with Ring Attractors
In standard PPO, the policy is directly parameterized by a neural network that outputs action probabilities. When integrating ring attractors, we modify this structure to incorporate the spatial relationships between actions through the ring topology.
Ring Attractor Policy Network. For PPO integration, we adapt the DL ring attractor structure from Section 3.2. While the value-based methods use the network to process Q-values, in PPO we process policy logits. For a policy network with parameters $\theta$, the initial policy logits are computed as:
$Z(s)=\Phi_{\theta}(s)$ (22)
This parallels Section 3.2, where $Q(s,a)=\Phi_{\theta}(s)$ represents the output of the network for value-based methods. The difference is that $Z(s)$ represents policy logits rather than action-values.
These logits are then processed through the ring attractor using the same weighted connection structure. The input signal $V(s)$ uses $Z(s)$ as input and follows the same form as in Eq. (12):
$V(s)_{m,n}=\frac{1}{\tau}Z(s)^{T}w^{I\rightarrow H}_{m,n}=\frac{1}{\tau}Z(s)^{T}e^{-\frac{d(m,n)}{\lambda}}$ (23)
Similarly, the hidden state $U(v)$ maintains the same recurrent dynamics as in Eq. (12). The key difference occurs in the final output layer, where instead of scaling the output to represent Q-values, we apply a softmax function to directly obtain action probabilities:
$\pi(a|s)=\text{softmax}(\tanh((V(s)+U(v))))$ (24)
A.3.2 Policy Gradient Experimental Results
As shown in Figure 4, PPO [29] with ring attractors (PPO-RA) demonstrates significant performance improvements over standard PPO in the OpenAI Gym Super Mario Bros environment [15].
Figure 4: Performance comparison: PPO-RA vs standard PPO in OpenAI Gym Super Mario Bros environment [15] with discrete action space, demonstrating adaptability to complex, game-like scenarios. The plot shows PPO-RA (red) consistently outperforming standard PPO (blue), with shaded regions representing standard deviation across 10 averaged seeds.
Our experimental results reveal an average performance increase of $71.4\%$ over standard PPO, with a final performance improvement of $67.5\%$ and a maximum performance gain of $161.4\%$ during training. These results demonstrate that the spatial awareness provided by ring attractors is beneficial across different RL paradigms, not just value-based methods. The performance gains in PPO-RA can be attributed to more efficient policy updates due to the preservation of spatial relationships between actions, smoother policy gradients resulting from the structured representation of the action space, and enhanced exploration through the natural diffusion of policy preferences across similar actions in the ring.
PPO-RA converges faster and achieves higher rewards than standard PPO. This suggests that ring attractors provide a general enhancement mechanism applicable to multiple RL approaches, showcasing that their utility may extend beyond the value-based methods explored in previous sections.
A.4 Validating Ring Attractor Contributions Through Ablation Studies
A.4.1 Exogenous Ring Attractor Model Ablation Study
@@
Figure 5: Ablation study comparing BDQN variants in OpenAI Gym Super Mario Bros [15]. The plot shows cumulative reward over 1 million frames for three models: Standard BDQN [1]; BDQNRA-UA with RA and Uncertainty Awareness (UA) implementing both the ring attractor behavior policy from Section 3.1.2 and the uncertainty quantification model from Section 3.1.3; and BDQNRA-RM, applying the same concepts from BDQNRA-UA, but randomly distributing the action space across the ring in each experiment. Displaying mean episodic returns over 10 averaged seeds.
A.4.2 Deep Learning Ring Attractor Model Ablation Study
@@
Figure 6: Ablation study results comparing the performance of the full RNN-based ring attractor model against a version with the circular weight distribution removed. The graph illustrates a significant performance drop for the Ms Pacman and Chopper Command environments in the Atari 100K benchmark [5]. This emphasizes the role of the circular topology in encoding spatial information and enhancing learning. Displaying mean episodic returns over 10 averaged seeds.
@@
A.4.3 Deep Learning Ring Attractor Model Evolution
In this appendix section, we analyze model dynamics with both forward pass ($V(s)$) and hidden-to-hidden ($U(v)$) weights made trainable, rather than the standard approach of fixed forward pass connections, as presented in Section 3.2. As shown in Fig. 7, the forward-pass connections preserve the ring structure over training time, with strong distance-dependent decay patterns maintained throughout the learning process. This may indicate that the network naturally favors maintaining spatial topology for transmitting sensory information on a per-frame basis.
@@ A.4.3 Deep Learning Ring Attractor Model Evolution
The hidden-to-hidden connections, depicted in Fig. 8, exhibit markedly different behavior. These connections evolve beyond their initial ring structure, developing specialized patterns that enable the encoding of environment-specific relationships between neurons in the hidden space. This flexibility in hidden-layer connectivity supports the learning of complex action relationships while building upon the structured spatial representation from the forward pass.
These findings validate our standard implementation approach described in Section 3.2, where forward pass connections are fixed and only hidden-to-hidden weights remain trainable. The natural preservation of ring structure in trainable forward weights suggests that this topology is inherently beneficial for processing spatial information, while adaptable hidden weights enable the task-specific learning demonstrated in our experimental results, Section 4.2.